A deck of 52 playing cards has 4 aces. The probability of getting one ace less than zero

1
Expert's answer
2022-03-01T18:05:05-0500

The number of aces in a pack of 52 card is 4, so the probability that the card drawn is an ace is,

P=4/52

Now, since one card is been chosen, then the total number of aces left is 3 and the total number of cards left is 52−1=51.

The probability that the card drawn is an ace is,

P=3/51


As two cards are been chosen, then the total number of aces left is 2 and the total number of cards left is 51−1=50.

The probability that the card drawn is an ace is,

P=2/50


Again, since three cards are been chosen, then the total number of aces left is 1 and the total number of cards left is 50−1=49.

The probability that the card drawn is an ace is,

P=1/49

By multiplication law of probability, we have

p= ( (4/52) * (3/51) * (2/50) * (1/49) )


Therefore the required probability equals 1/270725